# Log-majorizations between quasi-geometric type means for matrices

### v2

@@ 6 Concluding remarks
associated to a function $f$ on $(0,\infty)$ is jointly operator convex on $\mathbb{M}_{n}^{++}\times\mathbb{M}_{n}^{++}$, $n\in\mathbb{N}$, if $f$ is operator convex, while it is jointly operator concave if $f$ is operator monotone. Thus it might be expected to obtain a general joint convexity theorem for $\mathrm{Tr}\,\mathcal{P}_{f}(B^{p},A^{p})^{1/p}$ for certain $p>0$, that is the convexity counterpart of Theorem 5.1 when $f$ is operator convex on $(0,\infty)$. However, this is not possible except for $p=1$, in the case that $\mathcal{P}_{f}$ is jointly operator convex. Indeed, when $f$ is a linear function $f(x)=1-\alpha+\alpha x$ ($0<\alpha<1$), the joint convexity of $\mathrm{Tr}\,\mathcal{P}_{f}(B^{p},A^{p})^{1/p}=\mathrm{Tr}\,\mathcal{A}_{\alpha,p}(A,B)$ restricts to $1\leq p\leq 2$ by Proposition 5.2(2). When $f(x)=x^{\alpha}$ ($1<\alpha\leq 2$), the joint convexity of $\mathrm{Tr}\,\mathcal{P}_{f}(B^{p},A^{p})^{1/p}=\mathrm{Tr}\,G_{\alpha,p}$ restricts to $p\leq 1$. The intersection is only $p=1$. It might be possible that $\mathrm{Tr}\,\mathcal{P}_{f}(B^{p},A^{p})^{1/p}$ is jointly convex for a certain $p<1$ under the constraint of $f(0^{+})=0$.
@@ Appendix A Equality case in norm inequality for $G_{\alpha,p}$
By [39, Theorem 3.1] and [28, Lemma 2.2] again this implies that $\lambda\bigl(G_{\alpha}(e^{tH}t,e^{tK})^{1/t}\bigr)=\lambda(e^{\alpha H+(1-\alpha)K})$ for all $t>0$, so that
@@
where $\mathrm{Tr}_{nl}:\mathbb{M}_{n}\otimes\mathbb{M}_{l}\otimes\mathbb{M}_{m}\to\mathbb{M}_{m}$ is the partial trace. Let $\tau_{0}:=(nl)^{-1}I_{nl}\in\mathbb{M}_{n}\otimes\mathbb{M}_{l}$. Then $\tau_{0}\otimes\mathrm{Tr}_{nl}$ is the conditional expectation from $\mathbb{M}_{n}\otimes\mathbb{M}_{l}\otimes\mathbb{M}_{m}$ onto $I_{nl}\otimes\mathbb{M}_{m}$ with respect to the trace, and it is well known (see, e.g., [51, Sec. 7]) that for every $Z\in\mathbb{M}_{n}\otimes\mathbb{M}_{l}\otimes\mathbb{M}_{m}$,
@@ Appendix C Monotonicity under restricted CPTP maps
for all $a,b\in[0,\infty)^{n}$, $n\in\mathbb{N}$, such that $(\mathrm{diag}(a),\mathrm{diag}(b))\in\mathrm{dom}\,Q_{\alpha}$. Define the quantum divergence of Renyi type defined for $A,B\in\mathbb{M}_{n}^{+}$, $n\in\mathbb{N}$, with $B\neq 0$ as in (5.7) by
@@
We consider two special classes of quantum channels (CPTP maps). If $\Phi:\mathbb{M}_{n}\to\mathbb{M}_{m}$ is a positive trace-preserving map such that the range of $\Phi$ is included in a commutative subalgebra of $\mathbb{M}_{m}$ is called a quantum-classical channel. In this case, there are orthogonal projections $(P_{i})_{i=1}^{k}$ in $\mathbb{M}_{m}$ such that $\sum_{i=1}^{k}P_{i}=I_{m}$ and the range of $\Phi$ is included in $\bigoplus_{i=1}^{k}\mathbb{C}P_{i}$. Then we have a POVM $(M_{i})_{i=1}^{k}$ on $\mathbb{C}^{n}$ (see (5.3)) such that $\Phi(A)=\sum_{i=1}^{k}(\mathrm{Tr}\,M_{i}A/\mathrm{Tr}\,P_{i})P_{i}$ for $A\in\mathbb{M}_{n}$. In this way, a quantum-classical channel is essentially a POVM. The other way around, a positive trace preserving map $\Phi:\mathbb{C}^{n}\to\mathbb{M}_{m}$ is called a classical-quantum channel. Those quantum-classical and classical-quantum channels are are automatically CPTP maps.

### v3

@@ 6 Concluding remarks
associated to a function $f$ on $(0,\infty)$ is jointly operator convex on $\mathbb{M}_{n}^{++}\times\mathbb{M}_{n}^{++}$, $n\in\mathbb{N}$, if $f$ is operator convex, while it is jointly operator concave if $f$ is operator monotone. Thus it might be expected to obtain a general joint convexity theorem for $\mathrm{Tr}\,\mathcal{P}_{f}(B^{p},A^{p})^{1/p}$ for certain $p>0$, that is the convexity counterpart of Theorem 5.1, when $f$ is operator convex on $(0,\infty)$. However, this is not possible except for $p=1$, in the generic case that $\mathcal{P}_{f}$ is jointly operator convex. Indeed, when $f$ is a linear function $f(x)=1-\alpha+\alpha x$ ($0<\alpha<1$), the joint convexity of $\mathrm{Tr}\,\mathcal{P}_{f}(B^{p},A^{p})^{1/p}=\mathrm{Tr}\,\mathcal{A}_{\alpha,p}(A,B)$ restricts to $1\leq p\leq 2$ by Proposition 5.2(2). When $f(x)=x^{\alpha}$ ($1<\alpha\leq 2$), the joint convexity of $\mathrm{Tr}\,\mathcal{P}_{f}(B^{p},A^{p})^{1/p}=\mathrm{Tr}\,G_{\alpha,p}$ restricts to $p\leq 1$ by Proposition 5.16(1). The intersection is only $p=1$. It might be possible that $\mathrm{Tr}\,\mathcal{P}_{f}(B^{p},A^{p})^{1/p}$ is jointly convex for a certain $p<1$ under the constraint of $f(0^{+})=0$.
(8) We have checked that all of $\mathcal{A}_{\alpha,p}$ for $0<\alpha<1$ and $G_{\alpha,p}$, $R_{\alpha,p}$, $LE_{\alpha}$ for $\alpha\in(0,\infty)\setminus\{1\}$, where $p>0$, are neither jointly operator concave nor jointly operator convex, except for $\mathcal{A}_{\alpha,1}$, $G_{\alpha,1}$ for $0<\alpha<1$ and $G_{\alpha,1}$ for $1<\alpha\leq 2$. Therefore, two variable operator/matrix means other than Kubo–Ando’s operator means or operator perspectives for operator convex functions on $(0,\infty)$ can hardly be jointly operator concave/convex. Indeed, this is the reason why we have treated the joint concavity/convexity of the trace functions $\mathrm{Tr}\,\mathcal{M}_{\alpha,p}$ in Section 5. Nevertheless, it might be interesting to consider whether or not $\mathcal{M}_{\alpha,p}$ are jointly concave/convex in other orderings such as the entrywise eigenvalue order or the weak majorization (see [35, Sec. 2.2]).
Declaration of competing interest
The author has no competing interests to disclose.
@@ Appendix A Equality case in norm inequality for $G_{\alpha,p}$
By [41, Theorem 3.1] and [30, Lemma 2.2] again this implies that $\lambda\bigl(G_{\alpha}(e^{tH}t,e^{tK})^{1/t}\bigr)=\lambda(e^{\alpha H+(1-\alpha)K})$ for all $t>0$, so that
@@
where $\mathrm{Tr}_{nl}:\mathbb{M}_{n}\otimes\mathbb{M}_{l}\otimes\mathbb{M}_{m}\to\mathbb{M}_{m}$ is the partial trace. Let $\tau_{0}:=(nl)^{-1}I_{nl}\in\mathbb{M}_{n}\otimes\mathbb{M}_{l}$. Then $\tau_{0}\otimes\mathrm{Tr}_{nl}$ is the conditional expectation from $\mathbb{M}_{n}\otimes\mathbb{M}_{l}\otimes\mathbb{M}_{m}$ onto $I_{nl}\otimes\mathbb{M}_{m}$ with respect to the trace, and it is well known (see, e.g., [54, Sec. 7]) that for every $Z\in\mathbb{M}_{n}\otimes\mathbb{M}_{l}\otimes\mathbb{M}_{m}$,
@@ Appendix C Monotonicity under restricted CPTP maps
for all $a,b\in[0,\infty)^{n}$, $n\in\mathbb{N}$, such that $(\mathrm{diag}(a),\mathrm{diag}(b))\in\mathrm{dom}\,Q_{\alpha}$. Define the quantum divergence of Renyi type for $A,B\in\mathbb{M}_{n}^{+}$, $n\in\mathbb{N}$, with $B\neq 0$ as in (5.7) by
@@
We consider two special classes of quantum channels (CPTP maps). If $\Phi:\mathbb{M}_{n}\to\mathbb{M}_{m}$ is a positive trace-preserving map such that the range of $\Phi$ is included in a commutative subalgebra of $\mathbb{M}_{m}$ is called a quantum-classical channel. In this case, there are orthogonal projections $(P_{i})_{i=1}^{k}$ in $\mathbb{M}_{m}$ such that $\sum_{i=1}^{k}P_{i}=I_{m}$ and the range of $\Phi$ is included in $\bigoplus_{i=1}^{k}\mathbb{C}P_{i}$. Then we have a POVM $(M_{i})_{i=1}^{k}$ on $\mathbb{C}^{n}$ (see (5.3)) such that $\Phi(A)=\sum_{i=1}^{k}(\mathrm{Tr}\,M_{i}A/\mathrm{Tr}\,P_{i})P_{i}$ for $A\in\mathbb{M}_{n}$. In this way, a quantum-classical channel is essentially a POVM. The other way around, a positive trace-preserving map $\Phi:\mathbb{C}^{n}\to\mathbb{M}_{m}$ is called a classical-quantum channel. Those quantum-classical and classical-quantum channels are automatically CPTP maps.
@@
Data availability
No data was used for the research described in the article.